Show that when the intensity of a sound in decibel is triples the sound intensity in watt per metre is also triples. Find this sound level
1
Expert's answer
2019-05-27T11:02:54-0400
"\beta = 10 \log_{10}{\frac{I}{I_0}}"

"3\beta = 10 \log_{10}{\frac{3I}{I_0}}"

"3\beta = 10 \log_{10}{\frac{3I}{I_0}}+10 \log_{10}{3}"

"3\beta = \beta +10 \log_{10}{3}"

"2\beta =10 \log_{10}{3}"

"\beta =5 \log_{10}{3}=2.4 \ dB"